OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation x - 4y = 4. slope = 4; y intercept at (0, -1) slope = 1/4; y intercept at (0, 1) slope = 1/4; y intercept at (0, -1) slope = -4; y intercept at (0, 1)

OpenStudy (alexwee123):

put this into slope-intercept form x - 4y = 4. -4y=-x+4 y=1/4 x-1 so your slope is 1/4 and y-int is at (0,-1)

OpenStudy (anonymous):

thank you